# Location–Scale Calibration for Generalized Posterior

### v2

@@ 5.2 Experiment
We illustrate the finite-sample behavior of the plug-in location–scale calibration in the random-intercept linear mixed model with Huber loss and Gaussian prior with $n$-dependent scale described in Section 5.1. We fix $G=100$, $n_{i}=5$ (so $n=500$), and set $p=1$. We generate covariates from $\mathcal{N}(0,1)$ and contaminate Gaussian errors to induce model misspecification. Since no closed form is available for $\beta^{\lambda}$ in this example, we approximate it numerically by computing the penalized estimating equation estimator on large simulated datasets with $G=5,000$ and averaging over $1,000$ replications; see, e.g., oh2013simulated. For each learning rate $\eta$ on a logarithmic grid over $[0.01,100]$ we compare three procedures: (i) a frequentist Huber $M$-estimator with ridge penalty and sandwich-based Wald intervals; (ii) the generalized Bayes posterior based on $M_{n}$ and the corresponding $n$-dependent Gaussian prior; and (iii) the location–scale calibration applied to the same posterior draws, using the MAP estimate as an admissible center. For each $\eta$ we use $200$ Monte Carlo replications and record, for all three methods, the empirical coverage, the mean interval width, the mean bias, and the standard deviation of that bias. Further details of the data-generating mechanism, the numerical approximation of $\beta^{\lambda}$, and the Markov chain Monte Carlo settings are provided in Section E. The Python code for reproducing the experiments is available at https://github.com/shutech2001/ls-calib-gp.
@@ A.1 MAP centering
We first show that a maximum a posteriori estimator is an admissible center. Write the generalized log-posterior, up to an additive constant, as
@@
$\bigl\{\boldsymbol{J}_{n}\bigl(\bar{\boldsymbol{\theta}}_{n}\bigr)+\lambda_{n}\boldsymbol{H}_{\rho}\bigl(\tilde{\boldsymbol{\theta}}_{n}\bigr)\bigr\}\bigl(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{MAP}}-\check{\boldsymbol{\theta}}_{n}\bigr)=o_{p}\bigl(s_{n}^{-1/2}\bigr),$ (5)
for some random intermediate points $\bar{\boldsymbol{\theta}}_{n},\tilde{\boldsymbol{\theta}}_{n}$ on the line segment joining $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{MAP}}$ and $\check{\boldsymbol{\theta}}_{n}$. By Assumption 2.2 (iii), Assumption 2.1, and the consistency of $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{MAP}}$ and $\check{\boldsymbol{\theta}}_{n}$, we have
$\boldsymbol{J}_{n}\bigl(\bar{\boldsymbol{\theta}}_{n}\bigr)+\lambda_{n}\boldsymbol{H}_{\rho}\bigl(\tilde{\boldsymbol{\theta}}_{n}\bigr)\to_{p}\boldsymbol{J}_{\lambda}^{\star},$
and $\boldsymbol{J}_{\lambda}^{\star}$ is nonsingular. Hence the smallest eigenvalue of $\boldsymbol{J}_{n}(\bar{\boldsymbol{\theta}}_{n})+\lambda_{n}\boldsymbol{H}_{\rho}(\tilde{\boldsymbol{\theta}}_{n})$ is bounded away from zero in probability and its inverse is $O_{p}(1)$. Multiplying (5) by this inverse gives
@@ A.2 Generalized posterior mean centering
We next consider the generalized Bayes posterior mean $\boldsymbol{\theta}_{\mathrm{GB}}:=\mathbb{E}_{\Pi_{n}^{\eta}}[\boldsymbol{\theta}]$. To exploit Proposition 3.1 we center at the baseline estimator $\check{\boldsymbol{\theta}}_{n}$, which is itself an admissible center by Definition 2.5. Let
$\boldsymbol{Z}_{n}:=\sqrt{s_{n}}\bigl(\boldsymbol{\theta}-\check{\boldsymbol{\theta}}_{n}\bigr)$
@@
$\int_{\mathbb{R}^{p}}\bigl|q_{n}^{\check{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})-\mathcal{N}\bigl(\boldsymbol{x}\mid\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr)\bigr|\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow 0,\quad\boldsymbol{H}_{0}:=\eta\boldsymbol{J}_{\lambda}^{\star}.$
@@
$\boldsymbol{\delta}_{n}:=\sqrt{s_{n}}\bigl(\boldsymbol{\theta}_{\mathrm{GB}}-\check{\boldsymbol{\theta}}_{n}\bigr)=\int_{\mathbb{R}^{p}}\boldsymbol{x}q_{n}^{\check{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})\mathop{}\!\mathrm{d}\boldsymbol{x}=\mathbb{E}_{\Pi_{n}^{\eta}}\!\left[\boldsymbol{Z}_{n}\right].$
The total-variation convergence together with the fact that the second moments of $\boldsymbol{Z}_{n}$ converge to those of $\mathcal{N}(\boldsymbol{0},\boldsymbol{H}_{0}^{-1})$ (again by Proposition 3.1) implies $\boldsymbol{\delta}_{n}\to_{p}\boldsymbol{0}$. Thus
$\bigl\|\boldsymbol{\theta}_{\mathrm{GB}}-\check{\boldsymbol{\theta}}_{n}\bigr\|=\frac{\|\boldsymbol{\delta}_{n}\|}{\sqrt{s_{n}}}=o_{p}\bigl(s_{n}^{-1/2}\bigr).$ (6)
@@ A.3 One-step Newton centering
$\sqrt{s_{n}}\bigl\{\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\bigr\}\longrightarrow_{d}\mathcal{N}(\boldsymbol{0},\boldsymbol{K}^{\star}).$
@@
Hence, if in addition $\sqrt{s_{n}}(\lambda_{n}-\lambda)\to 0$ (e.g. when $\lambda_{n}\equiv\lambda$), we obtain
$\sqrt{s_{n}}\boldsymbol{F}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\longrightarrow_{d}\mathcal{N}(\boldsymbol{0},\boldsymbol{K}^{\star}).$
@@
$\sqrt{s_{n}}\bigl(\check{\boldsymbol{\theta}}_{n}-\boldsymbol{\theta}^{\lambda}\bigr)=-\boldsymbol{J}_{\lambda}^{\star-1}\sqrt{s_{n}}\boldsymbol{F}_{n}(\boldsymbol{\theta}^{\lambda})+o_{p}(1).$ (9)
@@
$\tilde{\boldsymbol{\theta}}_{n}^{(1)}-\boldsymbol{\theta}^{\lambda}$ $=\boldsymbol{\theta}_{\mathrm{GB}}-\boldsymbol{\theta}^{\lambda}-\boldsymbol{A}_{n}(\boldsymbol{\theta}_{\mathrm{GB}})^{-1}\boldsymbol{F}_{n}(\boldsymbol{\theta}_{\mathrm{GB}})$
$=\boldsymbol{\theta}_{\mathrm{GB}}-\boldsymbol{\theta}^{\lambda}-\boldsymbol{J}_{\lambda}^{\star-1}\bigl\{\boldsymbol{F}_{n}(\boldsymbol{\theta}^{\lambda})+\boldsymbol{A}_{n}(\boldsymbol{\theta}^{\lambda})(\boldsymbol{\theta}_{\mathrm{GB}}-\boldsymbol{\theta}^{\lambda})\bigr\}+o_{p}\bigl(s_{n}^{-1/2}\bigr)$
$=-\boldsymbol{J}_{\lambda}^{\star-1}\boldsymbol{F}_{n}(\boldsymbol{\theta}^{\lambda})+o_{p}\bigl(s_{n}^{-1/2}\bigr).$
@@
$\sqrt{s_{n}}\bigl(\tilde{\boldsymbol{\theta}}_{n}^{(1)}-\check{\boldsymbol{\theta}}_{n}\bigr)=\sqrt{s_{n}}\bigl(\tilde{\boldsymbol{\theta}}_{n}^{(1)}-\boldsymbol{\theta}^{\lambda}\bigr)-\sqrt{s_{n}}\bigl(\check{\boldsymbol{\theta}}_{n}-\boldsymbol{\theta}^{\lambda}\bigr)\longrightarrow_{p}\boldsymbol{0},$
@@ Appendix B Proofs of the main results
$\boldsymbol{u}=\sqrt{s_{n}}\bigl(\boldsymbol{\theta}-\check{\boldsymbol{\theta}}_{n}\bigr),\quad\boldsymbol{\theta}=\check{\boldsymbol{\theta}}_{n}+\frac{\boldsymbol{u}}{\sqrt{s_{n}}},$
@@
$M_{n}\!\left(\check{\boldsymbol{\theta}}_{n}+\frac{\boldsymbol{u}}{\sqrt{s_{n}}}\right)=M_{n}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)+\sqrt{s_{n}}\boldsymbol{U}_{n}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)^{\top}\boldsymbol{u}+\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{J}_{n}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)\boldsymbol{u}+R_{n,1}(\boldsymbol{u}),$ (11)
@@
$s_{n}\rho\left(\check{\boldsymbol{\theta}}_{n}+\frac{\boldsymbol{u}}{\sqrt{s_{n}}}\right)=s_{n}\rho\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)+\sqrt{s_{n}}\nabla\rho\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)^{\top}\boldsymbol{u}+\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{H}_{\rho}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)\boldsymbol{u}+R_{n,2}(\boldsymbol{u}),$ (12)
@@
For the prior remainder $r_{n}$, Assumption 2.3 states that $r_{n}$ is locally Lipschitz on $\mathcal{U}_{n}$ with Lipschitz constant $L_{n}=o_{p}(\sqrt{s_{n}})$. Lebourg’s mean value theorem for locally Lipschitz functions (see, e.g., clarke1990optimization) applied along the segment joining $\check{\boldsymbol{\theta}}_{n}$ and $\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}/\sqrt{s_{n}}$ gives, for each fixed $M<\infty$ and all $\|\boldsymbol{u}\|\leq M$,
$r_{n}\left(\check{\boldsymbol{\theta}}_{n}+\frac{\boldsymbol{u}}{\sqrt{s_{n}}}\right)=r_{n}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)+\frac{1}{\sqrt{s_{n}}}\xi_{n}(\boldsymbol{u})^{\top}\boldsymbol{u},$ (13)
@@
Substituting (11), (12) and (13) into $\log\pi_{n}^{\eta}$ at $\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}/\sqrt{s_{n}}$ yields
$\log\pi_{n}^{\eta}\!\left(\check{\boldsymbol{\theta}}_{n}+\frac{\boldsymbol{u}}{\sqrt{s_{n}}}\right)=C_{n}-\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{H}_{0,n}\boldsymbol{u}+R_{n}(\boldsymbol{u}),$
@@
The linear term in $\boldsymbol{u}$ arising from $M_{n}$ and $\rho$ are of order $\sqrt{s_{n}}$. The admissibility of $\check{\boldsymbol{\theta}}_{n}$ and the approximate penalized estimating equation imply
@@
$\eta\sqrt{s_{n}}\boldsymbol{U}_{n}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)-\eta\lambda_{n}\sqrt{s_{n}}\nabla\rho\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)=o_{p}(1)$
@@
$\left|\frac{1}{\sqrt{s_{n}}}\boldsymbol{\xi}_{n}(\boldsymbol{u})^{\top}\boldsymbol{u}\right|\leq\frac{L_{n}}{\sqrt{s_{n}}}M=o_{p}(1).$
@@
$\log\pi_{n}^{\eta}\!\left(\check{\boldsymbol{\theta}}_{n}+\frac{\boldsymbol{u}}{\sqrt{s_{n}}}\right)=C_{n}-\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{H}_{0,n}\boldsymbol{u}+R_{n}(\boldsymbol{u}),$
@@
Let $P_{n}^{0}$ denote the Gaussian measure $\mathcal{N}(\boldsymbol{0},\boldsymbol{H}_{0,n}^{-1})$ on $\mathbb{R}^{p}$, and let $\pi_{n}^{\circ}$ be the law of $\boldsymbol{u}$ induced by $\Pi_{n}^{\eta}$ via the above re-parametrization. The density of $\pi_{n}^{\circ}$ relative to $P_{n}^{0}$ is proportional to $\exp\{R_{n}(\boldsymbol{u})\}$. Since $\boldsymbol{H}_{0,n}\to_{p}\boldsymbol{H}_{0}\succ 0$, the eigenvalues of $\boldsymbol{H}_{0,n}$ are bounded away from zero and infinity with probability tending to one, and Gaussian tails under $P_{n}^{0}$ are uniformly controlled. Choosing $M<\infty$ large enough, we may assume $P_{n}^{0}(\|\boldsymbol{u}\|>M)$ is arbitrarily small uniformly in $n$. On $\{\|\boldsymbol{u}\|\leq M\}$, (14) implies $\sup_{\|\boldsymbol{u}\|\leq M}|R_{n}(\boldsymbol{u})|=o_{p}(1)$, and $\exp\{R_{n}(\boldsymbol{u})\}$ converges to $1$ uniformly on this set. A truncation and dominated convergence argument then yields
@@
$\int_{\mathbb{R}^{p}}\left|p_{n}^{\circ}(\boldsymbol{x})-\mathcal{N}\bigl(\boldsymbol{x}\mid\boldsymbol{0},\boldsymbol{H}_{0,n}^{-1}\bigr)\right|\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow_{p}0,$ (15)
where $p_{n}^{\circ}$ denotes the density of $\pi_{n}^{\circ}$. Since $\boldsymbol{H}_{0,n}\to_{p}\boldsymbol{H}_{0}$ and Gaussian laws depend continously on the covariance matrix in total variation when eigenvalues are uniformly bounded, we also have
$\int_{\mathbb{R}^{p}}\left|\mathcal{N}\bigl(\boldsymbol{x}\mid\boldsymbol{0},\boldsymbol{H}_{0,n}^{-1}\bigr)-\mathcal{N}\bigl(\boldsymbol{x}\mid\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr)\right|\mathop{}\!\mathrm{d}\boldsymbol{x}=o_{p}(1).$ (16)
@@
$\int_{\mathbb{R}^{p}}\left|q_{n}^{\check{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})-\mathcal{N}\left.\left(\boldsymbol{x}\right|\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\right)\right|\mathop{}\!\mathrm{d}\boldsymbol{x}\to_{p}0,\quad n\to\infty.$ (17)
where $q_{n}^{\check{\boldsymbol{\theta}}}$ is the density of $\sqrt{s_{n}}(\boldsymbol{\theta}-\check{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$.
@@
$\boldsymbol{\delta}_{n}:=\sqrt{s_{n}}\bigl(\tilde{\boldsymbol{\theta}}_{n}-\check{\boldsymbol{\theta}}_{n}\bigr).$
@@
$\|\boldsymbol{\delta}_{n}\|\leq\sqrt{s_{n}}\bigl\|\tilde{\boldsymbol{\theta}}_{n}-\boldsymbol{\theta}^{\lambda}\bigr\|+\sqrt{s_{n}}\bigl\|\check{\boldsymbol{\theta}}_{n}-\boldsymbol{\theta}^{\lambda}\bigr\|=o_{p}(1),$
and thus $\boldsymbol{\delta}_{n}\to_{p}\boldsymbol{0}$. Let $q_{n}^{\tilde{\boldsymbol{\theta}}_{n}}$ be the density of $\sqrt{s_{n}}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$. For each realization of the data,

### v3

@@ 5.2 Experiment
We illustrate the finite-sample behavior of the plug-in location–scale calibration in the random-intercept linear mixed model with Huber loss and Gaussian prior with $n$-dependent scale described in Section 5.1. We fix $G=100$ and $n_{i}=5$, so the total sample size is $n=500$, and set $p=1$. We generate covariates from $N(0,1)$ and contaminate Gaussian errors to induce model misspecification. Since no closed form is available for $\beta^{\lambda}$ in this example, we approximate it numerically by computing the penalized estimating equation estimator on large simulated datasets with $G=5,000$ and averaging over $1,000$ replications; see, e.g., oh2013simulated. For each learning rate $\eta$ on a logarithmic grid over $[0.01,100]$ we compare three procedures: (i) a frequentist Huber $M$-estimator with ridge penalty and sandwich-based Wald intervals; (ii) the generalized Bayes posterior based on $M_{n}$ and the corresponding $n$-dependent Gaussian prior; and (iii) the location–scale calibration applied to the same posterior draws, using the MAP estimate as an admissible center. For each $\eta$ we use $200$ Monte Carlo replications and record, for all three methods, the empirical coverage, the mean interval width, the mean bias, and the standard deviation of that bias. Further details of the data-generating mechanism, the numerical approximation of $\beta^{\lambda}$, and the Markov chain Monte Carlo settings are provided in Section E. The Python code for reproducing the experiments is available at https://github.com/shutech2001/ls-calib-gp.
@@ A.1 MAP centering
We first show that a maximum a posteriori (MAP) estimator is an admissible center. Write the generalized log-posterior, up to an additive constant, as
@@
$\bigl\{\boldsymbol{J}_{n}\bigl(\boldsymbol{\theta}^{\prime}_{n}\bigr)+\lambda_{n}\boldsymbol{H}_{\rho}\bigl(\boldsymbol{\theta}^{\prime\prime}_{n}\bigr)\bigr\}\bigl(\hat{\boldsymbol{\theta}}_{n}^{\mathrm{MAP}}-\check{\boldsymbol{\theta}}_{n}\bigr)=o_{p}\bigl(s_{n}^{-1/2}\bigr),$ (5)
for some random intermediate points $\boldsymbol{\theta}^{\prime}_{n},\boldsymbol{\theta}^{\prime\prime}_{n}$ on the line segment joining $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{MAP}}$ and $\check{\boldsymbol{\theta}}_{n}$. By Assumption 2.2 (iii), Assumption 2.1, and the consistency of $\hat{\boldsymbol{\theta}}_{n}^{\mathrm{MAP}}$ and $\check{\boldsymbol{\theta}}_{n}$, we have
$\boldsymbol{J}_{n}\bigl(\boldsymbol{\theta}^{\prime}_{n}\bigr)+\lambda_{n}\boldsymbol{H}_{\rho}\bigl(\boldsymbol{\theta}^{\prime\prime}_{n}\bigr)\to_{p}\boldsymbol{J}_{\lambda}^{\star},$
and $\boldsymbol{J}_{\lambda}^{\star}$ is nonsingular. Hence the smallest eigenvalue of $\boldsymbol{J}_{n}(\boldsymbol{\theta}^{\prime}_{n})+\lambda_{n}\boldsymbol{H}_{\rho}(\boldsymbol{\theta}^{\prime\prime}_{n})$ is bounded away from zero in probability and its inverse is $O_{p}(1)$. Multiplying (5) by this inverse gives
@@ A.2 Generalized posterior mean centering
We next consider the generalized Bayes posterior mean $\boldsymbol{\theta}_{\mathrm{GB}}:=E_{\Pi_{n}^{\eta}}(\boldsymbol{\theta})$. To exploit Proposition 3.1 we center at the baseline estimator $\check{\boldsymbol{\theta}}_{n}$, which is itself an admissible center by Definition 2.5. Let
$\boldsymbol{Z}_{n}:=s_{n}^{1/2}\bigl(\boldsymbol{\theta}-\check{\boldsymbol{\theta}}_{n}\bigr)$
@@
$\int_{\mathbb{R}^{p}}\bigl|q_{n}^{\check{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})-N\bigl(\boldsymbol{x}\mid\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr)\bigr|\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow 0,\quad\boldsymbol{H}_{0}:=\eta\boldsymbol{J}_{\lambda}^{\star}.$
@@
$\boldsymbol{\delta}_{n}:=s_{n}^{1/2}\bigl(\boldsymbol{\theta}_{\mathrm{GB}}-\check{\boldsymbol{\theta}}_{n}\bigr)=\int_{\mathbb{R}^{p}}\boldsymbol{x}q_{n}^{\check{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})\mathop{}\!\mathrm{d}\boldsymbol{x}=\mathbb{E}_{\Pi_{n}^{\eta}}\!\left[\boldsymbol{Z}_{n}\right].$
The total-variation convergence together with the fact that the second moments of $\boldsymbol{Z}_{n}$ converge to those of $N(\boldsymbol{0},\boldsymbol{H}_{0}^{-1})$, as established by Proposition 3.1, implies $\boldsymbol{\delta}_{n}\to_{p}\boldsymbol{0}$. Thus
$\bigl\|\boldsymbol{\theta}_{\mathrm{GB}}-\check{\boldsymbol{\theta}}_{n}\bigr\|=\|\boldsymbol{\delta}_{n}\|s_{n}^{-1/2}=o_{p}\bigl(s_{n}^{-1/2}\bigr).$ (6)
@@ A.3 One-step Newton centering
$s_{n}^{1/2}\bigl\{\boldsymbol{U}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)-\Psi\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\bigr\}\longrightarrow_{d}N(\boldsymbol{0},\boldsymbol{K}^{\star}).$
@@
Hence, if in addition $s_{n}^{1/2}(\lambda_{n}-\lambda)\to 0$, for example when $\lambda_{n}\equiv\lambda$, we obtain
$s_{n}^{1/2}\boldsymbol{F}_{n}\bigl(\boldsymbol{\theta}^{\lambda}\bigr)\longrightarrow_{d}N(\boldsymbol{0},\boldsymbol{K}^{\star}).$
@@
$s_{n}^{1/2}\bigl(\check{\boldsymbol{\theta}}_{n}-\boldsymbol{\theta}^{\lambda}\bigr)=-\boldsymbol{J}_{\lambda}^{\star-1}s_{n}^{1/2}\boldsymbol{F}_{n}(\boldsymbol{\theta}^{\lambda})+o_{p}(1).$ (9)
@@
$\begin{split}\tilde{\boldsymbol{\theta}}_{n}^{(1)}-\boldsymbol{\theta}^{\lambda}&=\boldsymbol{\theta}_{\mathrm{GB}}-\boldsymbol{\theta}^{\lambda}-\boldsymbol{A}_{n}(\boldsymbol{\theta}_{\mathrm{GB}})^{-1}\boldsymbol{F}_{n}(\boldsymbol{\theta}_{\mathrm{GB}})\\ &=\boldsymbol{\theta}_{\mathrm{GB}}-\boldsymbol{\theta}^{\lambda}-\boldsymbol{J}_{\lambda}^{\star-1}\bigl\{\boldsymbol{F}_{n}(\boldsymbol{\theta}^{\lambda})+\boldsymbol{A}_{n}(\boldsymbol{\theta}^{\lambda})(\boldsymbol{\theta}_{\mathrm{GB}}-\boldsymbol{\theta}^{\lambda})\bigr\}+o_{p}\bigl(s_{n}^{-1/2}\bigr)\\ &=-\boldsymbol{J}_{\lambda}^{\star-1}\boldsymbol{F}_{n}(\boldsymbol{\theta}^{\lambda})+o_{p}\bigl(s_{n}^{-1/2}\bigr).\end{split}$
@@
$s_{n}^{1/2}\bigl(\tilde{\boldsymbol{\theta}}_{n}^{(1)}-\check{\boldsymbol{\theta}}_{n}\bigr)=s_{n}^{1/2}\bigl(\tilde{\boldsymbol{\theta}}_{n}^{(1)}-\boldsymbol{\theta}^{\lambda}\bigr)-s_{n}^{1/2}\bigl(\check{\boldsymbol{\theta}}_{n}-\boldsymbol{\theta}^{\lambda}\bigr)\longrightarrow_{p}\boldsymbol{0},$
@@ Appendix B Proofs of the main results
$\boldsymbol{u}=s_{n}^{1/2}\bigl(\boldsymbol{\theta}-\check{\boldsymbol{\theta}}_{n}\bigr),\quad\boldsymbol{\theta}=\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}s_{n}^{-1/2},$
@@
$M_{n}\!\left(\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}s_{n}^{-1/2}\right)=M_{n}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)+s_{n}^{1/2}\boldsymbol{U}_{n}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)^{\top}\boldsymbol{u}+\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{J}_{n}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)\boldsymbol{u}+R_{n,1}(\boldsymbol{u}),$ (11)
@@
$s_{n}\rho\left(\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}s_{n}^{-1/2}\right)=s_{n}\rho\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)+s_{n}^{1/2}\nabla\rho\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)^{\top}\boldsymbol{u}+\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{H}_{\rho}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)\boldsymbol{u}+R_{n,2}(\boldsymbol{u}),$ (12)
@@
For the prior remainder $r_{n}$, Assumption 2.3 states that $r_{n}$ is locally Lipschitz on $\mathcal{U}_{n}$ with Lipschitz constant $L_{n}=o_{p}(s_{n}^{1/2})$. Lebourg’s mean value theorem for locally Lipschitz functions, as stated by clarke1990optimization, applied along the segment joining $\check{\boldsymbol{\theta}}_{n}$ and $\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}s_{n}^{-1/2}$ gives, for each fixed $M<\infty$ and all $\|\boldsymbol{u}\|\leq M$,
$r_{n}\left(\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}s_{n}^{-1/2}\right)=r_{n}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)+s_{n}^{-1/2}\xi_{n}(\boldsymbol{u})^{\top}\boldsymbol{u},$ (13)
@@
Substituting (11), (12) and (13) into $\log\pi_{n}^{\eta}$ at $\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}s_{n}^{-1/2}$ yields
$\log\pi_{n}^{\eta}\!\left(\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}s_{n}^{-1/2}\right)=C_{n}-\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{H}_{0,n}\boldsymbol{u}+R_{n}(\boldsymbol{u}),$
@@
The linear term in $\boldsymbol{u}$ arising from $M_{n}$ and $\rho$ are of order $s_{n}^{1/2}$. The admissibility of $\check{\boldsymbol{\theta}}_{n}$ and the approximate penalized estimating equation imply
@@
$\eta s_{n}^{1/2}\boldsymbol{U}_{n}\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)-\eta\lambda_{n}s_{n}^{1/2}\nabla\rho\bigl(\check{\boldsymbol{\theta}}_{n}\bigr)=o_{p}(1)$
@@
$\left|s_{n}^{-1/2}\boldsymbol{\xi}_{n}(\boldsymbol{u})^{\top}\boldsymbol{u}\right|\leq L_{n}s_{n}^{-1/2}M=o_{p}(1).$
@@
$\log\pi_{n}^{\eta}\!\left(\check{\boldsymbol{\theta}}_{n}+\boldsymbol{u}s_{n}^{-1/2}\right)=C_{n}-\frac{1}{2}\boldsymbol{u}^{\top}\boldsymbol{H}_{0,n}\boldsymbol{u}+R_{n}(\boldsymbol{u}),$
@@
Let $P_{n}^{0}$ denote the Gaussian measure $N(\boldsymbol{0},\boldsymbol{H}_{0,n}^{-1})$ on $\mathbb{R}^{p}$, and let $\pi_{n}^{\circ}$ be the law of $\boldsymbol{u}$ induced by $\Pi_{n}^{\eta}$ via the above re-parametrization. The density of $\pi_{n}^{\circ}$ relative to $P_{n}^{0}$ is proportional to $\exp\{R_{n}(\boldsymbol{u})\}$. Since $\boldsymbol{H}_{0,n}\to_{p}\boldsymbol{H}_{0}\succ 0$, the eigenvalues of $\boldsymbol{H}_{0,n}$ are bounded away from zero and infinity with probability tending to one, and Gaussian tails under $P_{n}^{0}$ are uniformly controlled. Choosing $M<\infty$ large enough, we may assume $P_{n}^{0}(\|\boldsymbol{u}\|>M)$ is arbitrarily small uniformly in $n$. On $\{\|\boldsymbol{u}\|\leq M\}$, (14) implies $\sup_{\|\boldsymbol{u}\|\leq M}|R_{n}(\boldsymbol{u})|=o_{p}(1)$, and $\exp\{R_{n}(\boldsymbol{u})\}$ converges to $1$ uniformly on this set. A truncation and dominated convergence argument then yields
@@
$\int_{\mathbb{R}^{p}}\left|p_{n}^{\circ}(\boldsymbol{x})-N\bigl(\boldsymbol{x}\mid\boldsymbol{0},\boldsymbol{H}_{0,n}^{-1}\bigr)\right|\mathop{}\!\mathrm{d}\boldsymbol{x}\longrightarrow_{p}0,$ (15)
where $p_{n}^{\circ}$ denotes the density of $\pi_{n}^{\circ}$. Since $\boldsymbol{H}_{0,n}\to_{p}\boldsymbol{H}_{0}$ and Gaussian laws depend continuously on the covariance matrix in total variation when eigenvalues are uniformly bounded, we also have
$\int_{\mathbb{R}^{p}}\left|N\bigl(\boldsymbol{x}\mid\boldsymbol{0},\boldsymbol{H}_{0,n}^{-1}\bigr)-N\bigl(\boldsymbol{x}\mid\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\bigr)\right|\mathop{}\!\mathrm{d}\boldsymbol{x}=o_{p}(1).$ (16)
@@
$\int_{\mathbb{R}^{p}}\left|q_{n}^{\check{\boldsymbol{\theta}}_{n}}(\boldsymbol{x})-N\left.\left(\boldsymbol{x}\right|\boldsymbol{0},\boldsymbol{H}_{0}^{-1}\right)\right|\mathop{}\!\mathrm{d}\boldsymbol{x}\to_{p}0,\quad n\to\infty.$ (17)
where $q_{n}^{\check{\boldsymbol{\theta}}}$ is the density of $s_{n}^{1/2}(\boldsymbol{\theta}-\check{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$.
@@
$\boldsymbol{\delta}_{n}:=s_{n}^{1/2}\bigl(\tilde{\boldsymbol{\theta}}_{n}-\check{\boldsymbol{\theta}}_{n}\bigr).$
@@
$\|\boldsymbol{\delta}_{n}\|\leq s_{n}^{1/2}\bigl\|\tilde{\boldsymbol{\theta}}_{n}-\boldsymbol{\theta}^{\lambda}\bigr\|+s_{n}^{1/2}\bigl\|\check{\boldsymbol{\theta}}_{n}-\boldsymbol{\theta}^{\lambda}\bigr\|=o_{p}(1),$
and thus $\boldsymbol{\delta}_{n}\to_{p}\boldsymbol{0}$. Let $q_{n}^{\tilde{\boldsymbol{\theta}}_{n}}$ be the density of $s_{n}^{1/2}(\boldsymbol{\theta}-\tilde{\boldsymbol{\theta}}_{n})$ under $\Pi_{n}^{\eta}$. For each realization of the data,